The table below shows the values of selected macroeconomics variables over a period:

YEAR 1 (R-Billion) YEAR 2 (R-Billion)



Investment 200 220







Saving 180 190



Export 100 110



Imports 120 140



Government Expenditure 150 160



Taxation 150 160



Equilibrium National Income 1 800 2 000







(a) Calculate the value of the multiplier for this economy.



(b) Should the full employment level of income be R-B2 255, by how much should government change its spending to reach this level of income during the next year given the multiplier has not changed.



(c) Evaluate whether or not this policy approach is effective in real life in achieving the desired level of GDP.
1
Expert's answer
2021-07-26T18:22:01-0400

(a) value of the multiplier .

Found by dividing equilibrium national income by government spending.

"=\frac {1800}{150}"

"=12"

The value of the multiplier is therefore "12\times".

(b)

Level of income=2255

Multiplier =12

"12=\frac{2255}{x}"

"12x=2255"

"x=\frac {2255}{12}"

"x=187.9"

The government should thus increase its spending by 27.9

"187.9-150=27.9"

(c)

This policy is effective in achieving the desired level of GDP. This is because it relates the equilibrium national level of income directly to the government spending.